A typical atomic nucleus is about 5.0 ร— 10โˆ’15๐‘š in radius. Use the

uncertainty principle to place a lower limit on the energy an electron must

have if it is to be part of a nucleus.[ans; 3.3 ร— 10โˆ’12๐ฝ]

II. A hydrogen atom is 5.3 ร— 10โˆ’11๐‘š in radius. Use the uncertainty principle


1
Expert's answer
2021-06-21T11:40:41-0400

"p=\frac{h}{4\pi r}=2.11\cdot 10^{-20}~\frac{kg\cdot m}{s},"

"E=c\sqrt{p^2+m^2c^2}\approx 3.3\cdot 10^{-12}~J."